Network Systems is introducing a new network card. Suppose Network Systems knows its fixed costs are $600,000, its variable costs are $500 per card, and it must sell 15,000 cards to break even the first year. What is the minimum price per unit it should charge?


1
Expert's answer
2021-10-05T17:07:42-0400

"15,000 =\frac{F}{P-V}"


"15,000 =\frac{ \$600,000}{\$500 - V}"


Cross multiply


"15,000(500 - V) =600,000"


"7,500,000 - 15,000V = 600,000"


"-15,000V = 600,000 -7,500,000"

"-15,000V = - 6,900,000"


"V = \frac{-6,900,000}{-15,000}"


"V=460"


The minimum price per unit it should charge is $460